### v2

# Gravity-aware Grasp Generation with Implicit Grasp Mode Selection for Underactuated Hands

Abstract
Learning-based grasp detectors typically assume a precision grasp, where each finger only has one contact point, and estimate the grasp probability. In this work, we propose a data generation and learning pipeline that can leverage power grasping, which has more contact points with an enveloping configuration and is robust against both positioning error and force disturbance. To train a grasp detector to prioritize power grasping while still keeping precision grasping as the secondary choice, we propose to train the network against the magnitude of disturbance in the gravity direction a grasp can resist (gravity-rejection score) rather than the binary classification of success. We also provide an efficient data generation pipeline for a dataset with gravity-rejection score annotation. In addition to thorough ablation studies, quantitative evaluation in both simulation and real-robot clarifies the significant improvement in our approach, especially when the objects are heavy.
I Introduction
Power grasp [1] is a grasping mode where the fingers and palm contact the object at multiple points. Since the object is enveloped by the fingers, the grasp is robust against both initial position error and post-grasp force disturbance. Current state-of-the-art machine-learning-based grasp detectors [2, 3, 4, 5, 6, 7, 8, 9], though, only assume precision grasp [1] where each finger has only one contact point. Even though precision grasping is easier to handle in both the data generation and learning aspects, its limited contact region makes the grasp fragile. A typical failure mode of a precision grasp by a parallel-jaw gripper is that the object rotates around the axis connecting the two contact points, during which translational displacement also occurs to result in dropping the object. The small contact region also imposes a high demand for accurate hand-eye calibration and hand positioning, which is challenging for low-cost manipulators or those under low-impedance control for safety. To this end, in this paper, we introduce a data generation and learning framework to learn power grasping in addition to the common precision grasping. Figure 2 (a) illustrates a typical precision grasp and power grasp.
Figure 1: While existing works only handle precision grasping, power grasping is more robust against both initial position error and post-grasp force disturbance. We propose a data generation pipeline and neural network model that prioritizes power grasping if applicable (left) but implicitly switches back to precision grasping if power grasping is not available due to collision with the table (middle) or other objects (right.)
Underactuated hands [10] have a larger number of joints than the number of actuation. This passive compliance allows the hand to adapt to various objects with non-typical geometries without explicitly controlling each joint. Since it makes power grasping easier, this work focuses on utilizing underactuated hands, or more specifically, Robotiq 2F-85 [11]. However, such mechanically automatic property makes learning power grasps challenging because whether the grasp is a power grasp or a precision grasp cannot be explicitly controlled but only decided through the interaction between the object and the hand’s link mechanism. In addition, since a power grasp requires the fingers to wrap around the object, it is not always possible due to collisions with the environment, such as the tabletop or other objects (see Fig. 1 middle and right.) To this end, we propose a neural network architecture that prioritizes power grasping if applicable but implicitly switches back to precision grasping if power grasping is not available. More specifically, while prior works [2, 3, 4, 5, 6, 7, 8, 9] only estimate the success probability of a grasp, we handle two values, namely (i) gravity-rejection score that represents the magnitude of disturbance in the gravity direction a grasp can support, which implicitly encourages power grasping since our data generation pipeline labels higher score for power grasps, and (ii) grasp validness which implicitly rejects grasps where the fingers will collide with the environment or it is too big for the hand to grasp.
Our contributions are as follows:
  • A grasp data generation pipeline for a dense annotation of grasp poses and the magnitude of disturbance in the gravity direction a grasp can support (gravity-rejection score) for both power grasping and precision grasping of underactuated hands.
  • A neural network model that prioritizes power grasping if applicable but implicitly switches back to precision grasping if power grasping is not available, by learning both gravity-rejection score and grasp validness.
  • A novel visualization approach to intuitively understand the network output. A thorough simulation evaluation and ablation study based on a novel evaluation procedure in which the grasp is performed against various object weights. A real-system quantitative evaluation that shows the significant improvement of our approach.
II Related Works
II-A Training Data for Learned Grasp Detectors
The quality of the training data plays a critical role in any machine-learning-based system. For grasp detectors, it is common to take the synthetic approach since gathering real physics data is costly. Mahler et al. [12] labeled grasps with a score based on the robust grasp wrench space analysis [13]. Fang et al. [14] and Wang et al. [7] adopted the same metrics as [12]. Fang et al. [15] further extended [7] by adding a score based on the distance from the object’s center of mass. Qin et al. [2] first sampled grasps for each object and then projected them to cluttered scenes to reduce the computational cost, with antipodal score as the metric. Zhao et al. [3] took the same approach as [2]. Cai et al. [9] also took the two-staged approach but employed Q1 metric [16] in addition to the antipodal score. In this paper, we also adopt the two-staged approach, but it differs from other works in the grasp generation of power grasps of an underactuated hand and evaluation metrics for the gravity-rejection score.
While analytical metrics are computationally efficient, it is difficult to capture all aspects of a contact, such as contact force distribution or motion after a slip. Kappler et al. [17] reported that a simulation-based approach achieves a better performance. Zhou et al. [18] and Eppner et al. [19] followed [17] in shaking the hand in a gravity-less simulation. Breyer et al. [5] evaluated grasps by actually lifting the object in a simulation and classified its success or failure. Jiang et al. [6] also followed [5]. Huang et al. [8] further shake the hand after lifting the object to select only robust grasps. Those prior works use the simulator to get a set of valid grasps. However, such binary classification suffers from the sim-to-real gap since the absolute value of success or failure highly depends on the simulator’s contact model and parameters. In this paper, we also take a simulation-based approach, but we acquire a continuous spectrum of the gravity-rejection score to train the neural network to prioritize grasps with higher safety margins, such as power grasps. In addition, we propose an approximation to allow the two-staged approach to relax the high computation cost of the simulation-based approaches. Section III details our data generation pipeline.
II-B Grasp Representation in Learned Graps Detectors
Figure 2: (a) Most works assume a precision grasp (top, bottom left) due to its simplicity. A power grasp (bottom right) allows a more robust grasp thanks to the larger number of contacts. (b) Illustration of Eq. 1 which approximate the multidimensional disturbance-rejection score to the gravity direction gravity-rejection score.
There are two major approaches to representing grasp with neural networks: sample-and-evaluate approach [20, 12, 21, 22, 3, 7] and feed-forward approach [23, 2, 4, 5, 6, 9, 8]. In the first approach, the network is only responsible for classifying and scoring the point cloud that is cropped and transformed to the hand coordinate; thus, it is extendable for power grasps. The multimodality that both power grasping and precision grasping may exist, though, prohibits using sample-efficient techniques such as sampling Darboux frame from the object surface [20, 22]. In order to avoid power grasping to induce a large computation time in the sampling stage due to the larger search space than other works, we opt for the feed-forward approach.
A large number of works in 3D grasp detection [22, 21, 2, 3, 4, 7, 8, 9], take point cloud as the input. Extending the combination of point cloud representation and feed-forward network to power grasps, though, is challenging. Qin et al. [2] regressed translation offset and rotation of the hand from each point on the object surface. Cai et al. [9] used the location and normal of the points on the object surface to decide five degrees of freedom of the hand and regressed the remaining in-plane rotation. Since [2, 9] interpreted the surface point as candidates that face the hand, a set of precision grasp and power grasp poses may be assigned to a single point, leading to representation ambiguity. Sundermeyer et al. [4] and Huang et al. [8], on the other hand, interpreted the surface points as candidates of contact and reported a superior performance. Contact points of a power grasp, however, significantly differ among the object size, making explicit contact representation challenging (see Fig. 2 (a).)
Morrison et al. [23] employed a 2D fully convolutional network to infer per-pixel grasp probability and hand yaw angle from a depth image. Breyer et al. [5] extended it to 3D grasp detection with TSDF (truncated signed distance function) as the input, where a 3D fully convolutional network estimated per-voxel grasp probability and rotation in SO(3). Jiang et al. [6] followed [5] and jointly learned an implicit representation. The fully convolutional architecture is suitable to be extended for power grasps since (i) a precision grasp and a power grasp are assigned to different voxels (see Fig. 2 (a) bottom, the distance from the object differs for the two grasp modes), (ii) a precision grasp and a power grasp that grasps similar location on the object tends to have a similar rotation, thus we can expect a continuous grasp rotation field. In this work, we use the same backbone as [5], but differ in that the network is trained against both the continuous gravity-rejection score and the binary classification. Section IV details our network implementation.
III Synthetic Data Generation with Gravity-Rejection Score
Given a set of scenes (in this work, 5K) with multiple objects (in this work, 1-5) placed in a cluttered way (see Fig. 3 bottom), our target is to generate a dense annotation (in this work, 1.2K on average for each scene) of grasps with both power grasp and precision grasp mode, and label the gravity-rejection score, which stands for the magnitude of disturbance in the gravity direction the grasp can support. Following the previous works [2, 3, 9], we first generate per-object grasp poses. Unlike the case of precision grasp, where grasp poses can be generated by sampling antipodal points on the object’s surface, it is not trivial to generate power grasps by a purely geometric approach. In addition, for underactuated hands, the final contact state is decided through the interaction between the object and the finger mechanism. We, therefore, take a simulation-based approach where we perturb a precision grasp and simulate the contact process by closing the hand against the object in a simulator. Through the contact, the object settles in the hand, and we record the hand’s pose relative to the object.
In order to train a network to prioritize power grasp, we need a grasp metric that takes a higher value for power grasp. While analytical metrics [16] are computationally efficient, they assume each contact is independent. For underactuated hands, though, the contacts are tightly coupled through the parallel link mechanism. Following [5, 6, 8], we also take a simulation approach for the grasp evaluation. Our approach, though, differs from theirs in two aspects: (i) instead of performing the simulation at the scene level, we perform at the object level and project the disturbance-rejection score to the gravity direction of the scenes, which drastically reduces the computation cost; (ii) instead of evaluating a grasp by a success/fail binary label, we treat them as a continuous value.
At the object level, for each grasp, we apply an increasing external force on the object until it leaves the hand, and record the magnitude of the force as disturbance-rejection score. By applying the force on the object’s center of mass, we can encourage the network to prioritize grasping near the center of mass, in addition to prioritizing power grasping. As the gravity direction is unknown at this stage, we perform this evaluation in multiple directions, resulting in a $n$-dimentional score where $n$ is the number of the directions. In this work, we selected $n=6$. Considering that we employed 75 objects and generated 3K grasps per object on average, we can reduce the number of simulations to 1/24 of the cases where each grasp is evaluated at the scene level by the simulation.
In order to project the $n$-dimentional disturbance-rejection score to the scene’s gravity direction to acquire the one-dimensional gravity-rejection score $f_{g}$, we take the following approximation:
$f_{g}=\underset{\{i|\bm{n}_{i}\cdot\bm{n}_{g}>\epsilon\}}{\min}\frac{f_{i}}{\bm{n}_{i}\cdot\bm{n}_{g}}$ (1)
where $\bm{n}_{g}\in\mathbb{R}^{3}$ is the unit vector heading the direction of gravity, $f_{i}$ is the $i$-th element of the disturbance-rejection score, $\bm{n}_{i}\in\mathbb{R}^{3}$ is the unit vector heading the direction of the $i$-th disturbance, and $\epsilon$ is a small positive value. Figure 2 (b) illustrates the approximation and Fig. 3 illustrates the whole data generation pipeline.
Figure 3: Schematic of our data generation pipeline. We sample precision grasps based on the antipodal score and coarse power grasps by heuristically shifting the precision grasps (top-left). Those initial grasps are refined through grasp simulation (top-right). We apply external force in multiple directions and acquire a multi-dimensional disturbance-rejection score for each grasp. Finally, we create multi-object scenes and project per-object grasp poses to the scene to acquire the training data (bottom). The disturbance-rejection score is projected to the scene’s gravity direction to acquire the gravity-rejection score.
IV Network Architecture and Training
Given a single depth image as the input, our target is to infer a set of grasp poses in SE(3) with gravity-rejection score. At the inference time, the robot tries to execute the first reachable and collision-free grasp with the highest gravity-rejection score. This differs from other works that estimate the probability of the grasp: a higher probability means the network is more confident, but it doesn’t mean the grasp is more robust; in our case the gravity-rejection score has a unit of [N], which directly represents the grasp’s robustness in the gravity direction.
As discussed in Section II-B, we take a 3D fully convolutional network as the backbone following [5], but with a gravity-rejection score regression head. Unlike the case of [5], where both positive and negative samples were included in the dataset, our data generation pipeline described in Sec. III only provides positive samples. We therefore train the gravity-rejection score head with L2 loss against positive examples only, leaving the invalid grasp regions as out-of-domain. While [5] introduced a heuristic filter to remove out-of-domain voxels, in this work, we propose to learn a grasp validness head to classify whether the voxel is out-of-domain. The evaluation in Sec. V-A shows that this classification head is also effective in collision avoidance or rejecting grasping an improper region. As the training data is densely annotated, we fill all voxels without a grasp label as invalid and use a weighted cross-entropy loss for the training.
V Experiment and Evaluation
V-A Analysis of Gravity-Rejection Score and Classification
[…]
Figure 5: Input TSDF volume (left) and output grasp volume (right) with $z=0.15$ cross-section. The markers with black-circle are voxels with positive classification. The short blue lines are the hand approach direction.
Figure 6: Input TSDF volume (left) and output grasp volume (right) with $z=0.08$ cross-section. The score prioritizes power grasps while the classification rejects grasps with collision or with too-big width.
Figure 7: Input TSDF volume (left) and output grasp volume (right) with $y=0$ cross-section. Even for the region invisible to the camera, the network well predicts grasp poses. The classification properly removes grasps where the palm may collide with other objects.
The color bar on the right plot of the figures stands for the gravity-rejection score, which is a continuous field that takes a higher value when it’s closer to the object. This distribution encourages power grasping, as shown in Fig. 2 (a). Since the gravity-rejection score head is only trained against valid grasps, its output for the out-of-domain voxels is incorrect, e.g., it takes an even higher value inside the object or still takes a medium score at too-far-from-the-object region. The markers surrounded by a black circle represent the voxel with a higher grasp validness than the threshold. The three figures show that this classification is highly effective in rejecting those too-close or too-far grasps. In addition, in Fig. 6, grasps at $x=0$, $y=\pm 0.05$ region are rejected. Indeed, grasping this region is impossible due to the limitation of the hand’s maximum openness of 85 mm; thus, training data does not contain such grasps. Similarly, in Fig. 7 grasps at $x=0.03$, $y=0.12$ are rejected. This is also reasonable since the hand’s palm will collide with the objects to grasp this concave part.
V-B Simulation Benchmark
Figure 8: Success ratio (SR) and clear ratio (CR) against multiple object weights. Our approach significantly outperformed the baselines, especially when the objects were heavy, showing the advantage of power grasp.
[…]
For the quantitative evaluation, we used the same metrics as [5, 2, 3, 9, 8, 6]: success ratio (SR) representing the number of successful grasps divided by the number of total trials, and clear ratio (CR) representing the number of successful grasps divided by the total number of objects. Each scene is terminated after two consecutive grasp failures or detection failures. We created 128 scenes with the mesh models provided by Breyer et al. [5]  [^3] We merged the two test mesh sets for packed and pile and removed those that didn’t fit our hand. which are used by [6, 8] also.
Even though some existing works evaluated their performance against various geoemetric properties (such as shape), there are limited examples that evaluate against various physical properties such as object weight or friction coefficient. In this work, we propose to evaluate the grasp against a varying object weight, since in the light-weight region, it captures whether the hand can reach the object, while in the heavy-weight region, it captures whether the hand can lift the object. Note that the capability to lift heavier objects is equivalent to the capability to lift the same objects with less grasp force, which is critical for handling fragile objects.
Figure 8 plots the SR and CR against the object weight from 0.1 kg to 1.5 kg. In the light-weight region (less than 0.5 kg), Ours-Precision performed similarly with VGN while VN-EdgeGraspNet outperformed both. This is reasonable because we share the same backbone as VGN, and the author of VN-EdgeGraspNet reported that their grasp representation is more advantageous. In the heavy-weight region (more than 1 kg), however, Ours-Precision outperformed VGN and performed on par with VN-EdgeGraspNet. This indicates that even limited to precision grasping, our gravity-rejection score is beneficial in grasping heavy objects. One hypothesis is that if the hand only supports precision grasping, introducing the gravigy-rejection score to the edge-grasp representation will provide a better score. Nevertheless, our focus is on leveraging power grasp. Ours, with all features including power grasp annotation, significantly outperformed the others, with 20.9 % improvement for both SR and CR in the case of 0.1 kg and 41.9 / 48.7 % improvement of SR / CR in the case of 1.5 kg compared with VN-EdgeGraspNet. Its superiority in the light-weight region indicates that power grasping is robust against positioning error, and that in the heavy-weight region indicates that it’s robust against gravity.
V-C Comparison on the Rotation Representation
Following Qin et al. [2], we also select rotation matrix as the rotation representation due to its continuity in the SO(3). However, [2] lacks a discussion on how to select the basis vectors of the rotation matrix, i.e., how we should define the TCP coordinate. A rotation matrix $R=[\bm{e}_{x},\bm{e}_{y},\bm{e}_{z}]\in\mathbb{R}^{3\times 3}$ is constructed by three basis vectors. The network estimates two of them $\tilde{\bm{e}}_{x},\tilde{\bm{e}}_{z}$ and reconstructs $R$ by:
$\mathbf{e}_{x}$ $=$ $\tilde{\mathbf{e}}_{x}/|\tilde{\mathbf{e}}_{x}|$
$\mathbf{e}_{z}$ $=$ $\mathbf{e}^{\prime}_{z}/|\mathbf{e}^{\prime}_{z}|,~{}\mathbf{e}^{\prime}_{z}=\tilde{\mathbf{e}}_{z}-(\mathbf{e}_{x}\cdot\tilde{\mathbf{e}}_{z})\mathbf{e}_{x}$ (2)
$\mathbf{e}_{y}$ $=$ $\mathbf{e}_{z}\times\mathbf{e}_{x}$
which means it prioritizes $\tilde{\bm{e}}_{x}$ to decide two dimensions of the rotation and supplementarily uses $\tilde{\bm{e}}_{z}$ to decide the remaining one, while doesn’t explicitly estimate the $\bm{e}_{y}$ direction. We name this order as $\bm{e}_{x}$-$\bm{e}_{z}$ and compare it with the opposite order $\bm{e}_{y}$-$\bm{e}_{z}$, in addition to a quaternion version (see Fig. 4 left for the TCP coordinate definition.) Table I summarizes the result. In addition to the large gap between the rotation matrix and quaternion, we also observe an improvement of $\bm{e}_{x}$-$\bm{e}_{z}$ over $\bm{e}_{y}$-$\bm{e}_{z}$. This indicates that $\bm{e}_{x}$ is easier for the network to learn than $\bm{e}_{y}$. In the other part of this paper, we learn $\bm{e}_{x}$ and $\bm{e}_{z}$, and use $\bm{e}_{x}$-$\bm{e}_{z}$ order for the inference.
TABLE I: Success ratio (SR) and clear ratio (CR) under different rotation representation
Rotation Matrix ($\mathbf{e}_{x}$-$\mathbf{e}_{z}$) Rotation Matrix ($\mathbf{e}_{y}$-$\mathbf{e}_{z}$) Quaternion
SR [%] 71.7 68.0 57.8
CR [%] 74.1 69.2 61.0
V-D Validation with a Physical System
In this subsection, we validate our approach with a physical system, as shown in Fig. 9. Since we only have physical access to the Robotiq hand, we trained a baseline version of our model to mimic [5] by (i) training data only contained precision grasps, (ii) we only used grasp validness and ignored gravity-rejection score for the inference, (iii) quaternion rotation representation. We selected nine kinds of objects from the YCB dataset [27], which we didn’t use for training, and created 20 scenes with 51 objects in total in a simulator. In the real system, we overlaid the scene’s mesh with the real-time-streamed point cloud and manually adjusted the objects’ pose in the real world. Depth image was acquired from a pair of RGB cameras and a learning-based stereo inference [28]. Table II summarizes the result, showing a significant improvement in our approach over the baseline. Note that we cannot compare the absolute SR / CR with the simulation evaluation in V-B due to the sim-to-real gap. Nevertheless, we can observe a consistent result that our approach outperforms the baselines in both environments.
Figure 9: Target objects (left) and capture (middle, right) of the validation with a physical system.
TABLE II: Success ratio (SR) and clear ratio (CR) in the real robot validation
Objects Trials Success SR [%] CR [%]
Ours 51 51 38 74.5 74.5
Baseline 51 48 31 64.6 60.8
VI Conclusion
In this paper, we presented a data generation and learning framework to leverage power grasping in addition to precision grasping. We proposed to train a neural network against the continuous gravity-rejection score, that is the magnitude of disturbance in the gravity direction a grasp can support, rather than the grasp probability, in order to let the network prioritize power grasping while still keeping precision grasping as a secondary choice. We also proposed a data generation pipeline to efficiently create a dataset with gravity-rejection score annotation. We provided a qualitative visualization and analysis of the network output and showed the effect of the combination of gravity-rejection score head and grasp validness head. For the quantitative evaluation, we proposed to evaluate the performance against varying object weights. A simulation evaluation proved a significant improvement in our approach, where power grasping improves the grasp robustness in both positioning accuracy error and gravity direction force disturbance. Finally, we provided real robot validation, which showed that our approach is also effective in a physical system.